# FLAME[Uncaptioned image]: Frozen Large Language Models Enable Data-Efficient Language-Image Pre-training

### v1

@@ 3.1 Framework Overview
where $f_{t}(y_{i},P_{k})\in\mathbb{R}^{d_{t}}$ generates text embeddings through the LLM with prompt $P_{k}$, $f_{v}(x_{i})\in\mathbb{R}^{d_{v}}$ extracts visual features through the encoder, $\phi(\cdot)$ is the two-layer MLP that projects visual features into the text embedding space, and $\tau$ is a learnable temperature parameter. The image-to-text loss $\mathcal{L}_{T}$ follows the same formulation with image and text roles reversed.
@@ 3.2 Multifaceted Prompt Distillation
While LLMs have shown promising capabilities in natural language understanding, their decoder-only architecture with causal attention inherently limits their ability to generate rich contextual representations. Previous works attempt to address this limitation either through bidirectional attention retraining [6] or by extracting the final token’s features [21, 55, 52]. However, these approaches only produce a single universal representation, which is insufficient for language-image pre-training where images contain multiple semantic facets from objects to scenes. To address this challenge, we introduce multifaceted prompt distillation, a systematic approach that decomposes visual semantics into a hierarchical structure mirroring human visual perception, from local object properties to global scene understanding.
@@
  • Scene Level: Abstracts high-level concepts such as atmosphere and emotion, capturing the holistic context that humans naturally perceive beyond individual objects and actions.
Figure 3: Facet-decoupled attention. By streamlining all prompts with a shared prefix and applying this facet-decoupled attention mask, the overhead of feature extraction is greatly reduced.
@@ 3.3 Facet-Decoupled Attention
Our key insight is that all prompts share an identical prefix structure which dominates the input sequence length. Specifically, each prompt contains the core context: “Detailed image description: $y_{i}$. After thinking step by step,”. This observation motivates us to transform $K$ independent sequences into a single structured input:
@@
(a) Text Retrieval
(b) Image Retrieval
Figure 4: Multilingual zero-shot retrieval recall@1 results on Crossmodal-3600. Left: text retrieval. Right: image retrieval. Despite being trained solely on English datasets, FLAME achieves outstanding average performance across all 36 languages, surpassing mSigLIP, which is trained on the multilingual WebLI dataset with 100 languages.
@@ 4.1 Datasets and Implementation Details
Dataset Details.
We use two primary datasets for language-image pre-training: CC3M [37] and YFCC15M [30]. Our ablation studies are primarily conducted using the CC3M dataset. Given the differences in the availability of web-crawled data, we re-implement the data synthesis pipeline from DreamLIP [54] using MiniCPM-Llama3-V 2.5111https://huggingface.co/openbmb/MiniCPM-Llama3-V-2_5 to generate both long and short captions. The raw captions are also included during training. All these datasets will be released to the community to benefit other research.
Implementation Details.
Our method is implemented using OpenCLIP [^2] https://github.com/mlfoundations/open_clip [18] and trained on a single node equipped with 8 NVIDIA A800 GPUs. The evaluation follows the CLIP benchmark protocol [^3] https://github.com/LAION-AI/CLIP_benchmark or the respective official benchmark protocols. We use ViT-B/16 [13, 5] as the default visual encoder and Mistral-Nemo [^4] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 [20] as the default large language model (LLM). For tasks with short text input, we apply a single prompt for inference, which is the first prompt at the scene level provided in the supplementary material. For a fair comparison with previous methods, we train all models for 32 epochs. Input images are resized to 224 $\times$ 224. The learnable temperature parameter $\tau$ is initialized to 0.07. More detailed hyper-parameters are provided in the supplementary material.
4.2 Zero-shot Image-Text Retrieval
@@
l—l—cc—cc & S4V-val Urban-1k
Method Dataset I2T T2I I2T T2I
CLIP [35] CC3M 21.4 20.2 10.7 9.6
Long-CLIP [53] CC3M+S4V 51.3 46.1 15.5 18.5
FLAME CC3M 85.6 80.0 65.3 66.6
CLIP [35] YFCC15M 57.1 45.9 30.4 23.6
Long-CLIP [53] YFCC15M+S4V 77.2 77.6 40.5 46.1
FLAME YFCC15M
94.1 93.2 84.0 87.9
CLIP [35] WIT-400M 78.2 79.6 67.5 53.3
{NiceTabular}
l—l—cccccc—cccccc Text Retrieval Image Retrieval
@@
FLAME excels in long-context image-text retrieval tasks, as demonstrated by evaluations on ShareGPT4V-val [9] and Urban-1k [53]. Despite being trained on YFCC15M, FLAME achieves an 87.9% recall@1 in Urban-1k text-to-image retrieval, representing a 34.6% improvement over CLIP trained on the much larger WIT-400M dataset. Additionally, FLAME demonstrates an average improvement of 14.8% on ShareGPT4V-val. We also re-implement the recent method based on positional encoding interpolation [53], i.e., pre-training a CLIP model on CC3M and YFCC15M, then using their official code [^5] https://github.com/beichenzbc/Long-CLIP for fine-tuning on ShareGPT4V. The gap in results further demonstrates the advantage of FLAME’s paradigm shift over the interpolation-based architectural change on the standard CLIP text encoder.
@@
FLAME leverages the intrinsic capabilities of pre-trained LLMs, enabling effective multilingual generalization. Although the model is trained on English datasets, it can be directly applied to downstream tasks in multiple languages without additional training. We evaluate this multilingual retrieval capability by conducting zero-shot retrieval across 36 languages from the Crossmodal-3600 dataset [40]. As shown in Figure 4, FLAME outperforms WIT-400M-trained CLIP [35] across all 36 languages. Specifically, FLAME achieves an average image-to-text recall@1 of 51.7% and text-to-image recall@1 of 49.0%, surpassing CLIP by margins of 44.4% and 44.5%, respectively. Notably, FLAME outperforms mSigLIP [51], which is trained on the multilingual WebLI [10] dataset with 100 languages, by 4.3% in image-to-text recall@1 and 14.6% in text-to-image recall@1.
@@
l—l—cccccccccc—c—c Dataset Method Food-101 CIFAR-10 CIFAR-100 SUN397 Cars Aircraft DTD Pets Caltech-101 Flowers Average ImageNet
@@
Table 4: Multilingual ImageNet1k classification. CLIP is trained on WIT-400M. CN-CLIP is pre-trained on WIT-400M and fine-tuned on 200M Chinese data. NLLB-CLIP is pre-trained on WIT-400M and fine-tuned on LAION-COCO-NLLB with 200 languages. FLAME is trained on YFCC15M. Unlike other methods, FLAME generalizes well from monolingual training to multilingual evaluation.
@@ 4.3 Image Classification
l—l—ccccccc—c Dataset Method Food-101 CIFAR-10 CIFAR-100 Cars Aircraft DTD Caltech-101 Average
@@
4.4 Ablation Studies
@@
c—cc—cc—cc—cc—cc ShareGPT4V-val Urban-1k MSCOCO Flickr30k Classfication
@@
Our results show that using too few prompts leads to suboptimal performance, particularly on the long-context benchmarks. This phenomenon is due to the incomplete extraction and learning of features from long captions. Notably, the performance drop is less pronounced on classification tasks involving simpler short text inputs. On the other hand, when more than 7 prompts are used, the accuracy improvement plateaus. This suggests that beyond a certain point, the additional prompts provide diminishing returns, and the feature extraction from the long caption reaches a saturation point where further prompts may even hinder performance on short-context and classification tasks. Therefore, we use 7 as the default number of prompts to achieve a more balanced performance.
@@
To assess the robustness and scalability of our framework, we conducted ablation studies using different backbone architectures. These included various text encoders (Mistral-7B [^6] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, Mistral-Nemo, and Llama-3.1-8B [^7] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct) and visual encoders (ViT-S/16, ViT-B/16, and ViT-L/14). Whether reducing or increasing the architecture size, our framework continues to perform robustly. For instance, when using Mistral-7B, our model achieves an ImageNet top-1 accuracy of 34.4% and an average accuracy of 39.6% on downstream datasets, both of which significantly surpass the prior state-of-the-art results (31.1% and 31.0%, respectively) [54]. Additionally, we observe that performance improvements scale with the size of the visual encoder. For instance, FLAME with the ViT-L/14 encoder increases ImageNet top-1 accuracy by 22.5% over the CLIP baseline, while using ViT-S/16 yields a 19.0% improvement, underscoring the scalability of our approach.
Figure 5: Embedding and training overhead. FDA denotes facet-decoupled attention. While online training is hindered by inefficiencies, our approach of freezing LLMs allows for offline embedding to lower the overhead. Our facet-decoupled attention mechanism further boosts the efficiency of offline embedding, resulting in a 3.6x speedup.
@@
We evaluate the embedding and training overhead on CC3M, with results illustrated in Figure 5. For the recent method [54], we utilize their official code [^8] https://github.com/zyf0619sjtu/DreamLIP to measure training time. As shown, online training suffers from inefficiencies due to the frequent forward passes of LLMs. Our option of freezing LLMs, however, enables offline embedding preprocessing, significantly reducing unnecessary overhead. Furthermore, our proposed facet-decoupled attention greatly enhances the efficiency of offline embedding, achieving a 3.6x speedup. Our training time remains competitive with CLIP and outperforms that of the recent method [54].

### v3

@@ 3.1 Framework Overview
where $f_{t}(y_{i},P_{k})\in\mathbb{R}^{d_{t}}$ generates text embeddings through the LLM with prompt $P_{k}$, $f_{v}(x_{i})\in\mathbb{R}^{d_{v}}$ extracts visual features through the encoder, $\phi(\cdot)$ is the MLP that projects visual features into the text embedding space, and $\tau$ is the learnable temperature. The image-to-text loss $\mathcal{L}_{T}$ follows the same formulation with image and text roles reversed.
@@ 3.2 Multifaceted Prompt Distillation
While LLMs have shown promising capabilities in natural language understanding, their decoder-only architecture with causal attention inherently limits their ability to generate rich contextual representations. Previous works attempt to address this limitation either through bidirectional attention retraining [6] or by extracting the final token’s features [22, 65, 61]. However, these approaches only produce a single universal representation, which is insufficient for language-image pre-training where images contain multiple semantic facets from objects to scenes. To address this challenge, we introduce multifaceted prompt distillation, a systematic approach that decomposes visual semantics into a hierarchical structure mirroring human visual perception, from local object properties to global scene understanding. This semantic decomposation aligns with well-established visual understanding paradigms [32, 29, 40].
@@
  • Scene Level: Abstracts high-level concepts such as atmosphere and emotion, capturing the holistic context humans naturally perceive beyond individual objects and actions.
Figure 3: Facet-decoupled attention. By streamlining all prompts with a shared prefix and applying this facet-decoupled attention mask, the overhead of feature extraction is greatly reduced. The positions in red indicate the features to be extracted.
@@ 3.3 Facet-Decoupled Attention
Our key insight is that all prompts share an identical prefix structure which dominates the input sequence length. Specifically, each prompt contains the core context: “Detailed image description: $y_{i}$. After thinking step by step,". This observation motivates us to transform $K$ independent sequences into a single structured input:
@@
{NiceTabular}
l|l|cc|cc|cc|cc & S4V-val Urban-1k DCI DOCCI-test
Method Dataset I2T T2I I2T T2I I2T T2I I2T T2I
CLIP [41] CC3M 21.4 20.2 10.7 9.6 8.3 7.5 8.6 7.0
Long-CLIP [62] CC3M+S4V 51.3 46.1 15.5 18.5 13.8 14.2 14.7 12.6
FLAME CC3M 85.6 80.0 65.3 66.6 50.8 49.3 54.5 51.9
CLIP [41] YFCC15M 57.1 45.9 30.4 23.6 22.0 19.1 26.5 23.3
Long-CLIP [62] YFCC15M+S4V 77.2 77.6 40.5 46.1 29.4 29.7 35.0 33.5
FLAME YFCC15M
94.1 93.2 84.0 87.9 66.1 68.1 75.8 76.2
CLIP [41] WIT-400M 78.2 79.6 67.5 53.3 45.4 43.0 60.7 57.0
@@ 4.1 Datasets and Implementation Details
Figure 4: Multilingual zero-shot retrieval recall@1 results on Crossmodal-3600 (text-to-image retrieval). Despite being trained solely on English datasets, FLAME achieves outstanding average performance across all 36 languages, surpassing mSigLIP, which is trained on the multilingual WebLI dataset with 100 languages. The image-to-text results are provided in the supplementary materials.
@@
l|l|cccccc|cccccc Text Retrieval Image Retrieval
@@
Dataset Details.
We use two primary datasets for language-image pre-training: CC3M [43] and YFCC15M [33]. Our ablation studies are primarily conducted using the CC3M dataset. Given the differences in the availability of web-crawled data, we re-implement the data synthesis pipeline from DreamLIP [63], using MiniCPM-Llama3-V 2.5 to generate both long and short captions. The prompts are aligned with DreamLIP’s approach, i.e., “Describe the image in detail/short." The raw captions are also included during training. All these datasets will be released to the community to benefit other research.
Implementation Details.
Our method is implemented using OpenCLIP [^1] https://github.com/mlfoundations/open_clip [19] and trained on a single node equipped with 8 NVIDIA A800 GPUs. The evaluation follows the CLIP benchmark protocol [^2] https://github.com/LAION-AI/CLIP_benchmark or the respective official benchmark protocols. We use ViT-B/16 [13, 41, 5] as the default visual encoder and Mistral-Nemo  [21] as the default large language model (LLM). For tasks with long text input, we simply calculate the average cosine similarity between an image embedding and multiple text embeddings, avoiding complex methods. For tasks with short text input, we apply a single prompt for inference, which is the first prompt at the scene level provided in the supplementary material. For a fair comparison with previous methods, we train all models for 32 epochs. Input images are resized to 224 $\times$ 224. The learnable temperature parameter $\tau$ is initialized to 0.07. More detailed hyper-parameters are provided in the supplementary material.
4.2 Zero-shot Image-Text Retrieval
@@
FLAME excels in long-context image-text retrieval tasks, as demonstrated by evaluations on ShareGPT4V-val [9] and Urban-1k [62]. Despite being trained on YFCC15M, FLAME achieves an 87.9% recall@1 in Urban-1k text-to-image retrieval, representing a 34.6% improvement over CLIP trained on the much larger WIT-400M dataset. Additionally, FLAME demonstrates an average improvement of 14.8% on ShareGPT4V-val. We also re-implement the recent method based on positional encoding interpolation [62], i.e., pre-training a CLIP model on CC3M and YFCC15M, then using their official code for fine-tuning on ShareGPT4V. The gap in results further demonstrates the advantage of FLAME’s paradigm shift over the interpolation-based architectural change on the standard CLIP text encoder.
@@
FLAME leverages the intrinsic capabilities of pre-trained LLMs, enabling effective multilingual generalization. Although the model is trained on English datasets, it can be directly applied to downstream tasks in multiple languages without additional training. We evaluate this multilingual retrieval capability by conducting zero-shot retrieval across 36 languages from the Crossmodal-3600 dataset [46]. As shown in Figure 4 and Figure C1 in the supplementary materials, FLAME outperforms WIT-400M-trained CLIP [41] across all 36 languages. Specifically, FLAME achieves an average image-to-text recall@1 of 51.7% and text-to-image recall@1 of 49.0%, surpassing CLIP by margins of 44.4% and 44.5%, respectively. Notably, FLAME outperforms mSigLIP [60], which is trained on the multilingual WebLI [10] dataset with 100 languages, by 4.3% in image-to-text recall@1 and 14.6% in text-to-image recall@1.
@@
l|l|cccccccccc|c|c Dataset Method Food-101 CIFAR-10 CIFAR-100 SUN397 Cars Aircraft DTD Pets Caltech-101 Flowers Average ImageNet
@@
Table 4: Multilingual ImageNet classification. CLIP is trained on WIT-400M. CN-CLIP is pre-trained on WIT-400M and fine-tuned on 200M Chinese data. NLLB-CLIP is pre-trained on WIT-400M and fine-tuned on LAION-COCO-NLLB with 200 languages. FLAME is trained on YFCC15M. Unlike other methods, FLAME generalizes from monolingual training to multilingual evaluation.
@@ 4.3 Image Classification
l|l|ccccccc|c Dataset Method Food-101 CIFAR-10 CIFAR-100 Cars Aircraft DTD Caltech-101 Average
@@
Figure 5: Semantic interpretability. Based on vocabulary mapping, FLAME achieves patch-to-word translation with competent interpretability of language-image alignment. We apply average pooling to reduce the number of words for a clearer presentation.
@@
c|cc|cc|cc|cc|cc ShareGPT4V-val Urban-1k MSCOCO Flickr30k Classfication
@@
4.4 Semantic Interpretability
Since we align the visual encoder with the last hidden layer of the LLM, we can input each visual output into the projection layer of the LLM to get the next-token logits. We then look up the token with the highest probability in the LLM’s vocabulary, allowing each image patch to be semantically interpreted as a word. The overall process can be expressed as: $C_{i}=M(h(f_{I}(x_{i})))$, where $h$ denotes the projection head of the LLM, $M$ represents the mapping process that identifies the word with the highest probability in the LLM’s vocabulary. As shown in Figure 5, the mapped words reflect meaningful semantic properties such as categories (e.g., cat, plate), attributes (e.g., circular), and numerical values (e.g., four). This semantic mapping enhances interpretability by translating image features into understandable text.
4.5 Ablation Studies
@@
Our results show that using too few prompts leads to suboptimal performance, particularly on the long-context benchmarks. This phenomenon is due to the incomplete extraction and learning of features from long captions. Notably, the performance drop is less pronounced on classification tasks involving simpler short text inputs. On the other hand, when more than 7 prompts are used, the accuracy improvement plateaus. Therefore, we use 7 as the default number of prompts to achieve a more balanced performance.
@@
To assess the robustness and scalability of our framework, we conducted ablation studies using different backbone architectures. These included various text encoders (Mistral-7B , Mistral-Nemo, and Llama-3.1-8B ) and visual encoders (ViT-S/16, ViT-B/16, and ViT-L/14). Whether reducing or increasing the architecture size, our framework continues to perform robustly. For instance, when using Mistral-7B, our model achieves an ImageNet top-1 accuracy of 34.4% and an average accuracy of 39.6% on downstream datasets, both of which significantly surpass the prior state-of-the-art results (31.1% and 31.0%, respectively) [63]. Additionally, we observe that performance improvements scale with the size of the visual encoder. For instance, FLAME with the ViT-L/14 encoder increases ImageNet top-1 accuracy by 22.5% over the CLIP baseline, while using ViT-S/16 yields a 19.0% improvement, underscoring the scalability of our approach.
Figure 6: Embedding and training overhead. FDA denotes facet-decoupled attention. While online training is hindered by inefficiencies, our approach of freezing LLMs allows for offline embedding to lower the overhead. Our FDA further boosts the efficiency of offline embedding, resulting in a 3.6x speedup.
@@
We evaluate the embedding and training overhead on CC3M, with results illustrated in Figure 6. For the recent method [63], we utilize their official code to measure training time. As shown, online training suffers from inefficiencies due to the frequent forward passes of LLMs. Our option of freezing LLMs, however, enables offline embedding preprocessing, significantly reducing unnecessary overhead. Furthermore, our proposed facet-decoupled attention greatly enhances the efficiency of offline embedding, achieving a 3.6x speedup. Our training time remains competitive with CLIP and outperforms that of the recent method [63].
@@ 5 Conclusion and Limitation
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (62206271), the Fundamental Research Funds for the Central Universities No.xxj032023020, the Shenzhen Key Technical Projects under Grant (JSGG20220831105801004, JCYJ20220818101406014, CJGJZD2022051714160501), the Guangdong University Featured Innovation Program Project (2024KTSCX026), and the CAAI-MindSpore Open Fund, developed on OpenI Community.
Supplementary Material
This document is structured as follows:
  • In Section A, we provide details of the hyper-parameters.
  • In Section B, we present the full list of multifaceted prompts used in our experiments.
  • In Section C, we conduct additional experiments, including: 1) multilingual image-to-text retrieval, 2) vision-language compositionality, 3) few-shot image comprehension, 4) ablations on individual semantic levels, and 5) preliminary exploration of scalability.
  • In Section D, we provide visualizations of text-to-image retrieval.
A Hyper-Parameters
Table A and A provide the pre-training hyperparameters used for CC3M and YFCC15M, respectively. These hyper-parameters are set similarly to those used in previous methods [14, 63] for fair comparisons.
{NiceTabular}
l|l Config Value
Batch size 4,096
Optimizer AdamW [37]
Learning rate 5e-4
Weight decay 0.5
Adam $\beta_{1}$, $\beta_{2}$ 0.9, 0.98
Adam $\epsilon$ 1e-8
Total epochs 32
Warm up iterations 2,000
Learning rate schedule cosine decay
Table A1: Hyper-parameters for CC3M.
{NiceTabular}
l|l Config Value
Batch size 8,192
Optimizer AdamW [37]
Learning rate 5e-4
Weight decay 0.2
Adam $\beta_{1}$, $\beta_{2}$ 0.9, 0.98
Adam $\epsilon$ 1e-8
Total epochs 32
Warm up iterations 2,000
Learning rate schedule cosine decay
Table A2: Hyper-parameters for YFCC15M.
B Full List of Multifaceted Prompts
We provide all the designed multifaceted prompts here. The distinct part in each prompt is marked. We also annotate the default prompts for long and short text input.
Entity Level:
  • Detailed image description: “$y_{i}$". After thinking step by step, the category of the main object in this image means in just one word:" (default for long input)
  • Detailed image description: “$y_{i}$". After thinking step by step, the prominent characteristic or pattern of the main object in this image means in just one word:" (default for long input)
  • Detailed image description: “$y_{i}$". After thinking step by step, the category of the minor object in this image means in just one word:" (default for long input)
  • Detailed image description: “$y_{i}$". After thinking step by step, the prominent characteristic or pattern of the minor object in this image means in just one word:" (default for long input)
Interaction Level:
  • Detailed image description: “$y_{i}$". After thinking step by step, the primary action or event taking place in this image means in just one word:" (default for long input)
  • Detailed image description: “$y_{i}$". After thinking step by step, the positioning layout or spatial relationship in this image means in just one word:"
Scene Level:
  • Detailed image description: “$y_{i}$". After thinking step by step, this image description means in just one word:" (default for long and short inputs)
  • Detailed image description: “$y_{i}$". After thinking step by step, the overall atmosphere or emotion of this image means in just one word:" (default for long input)
  • Detailed image description: “$y_{i}$". After thinking step by step, the dominant color or color combination of this image means in just one word:"
C Additional Experiments
Figure C1: Multilingual zero-shot image-to-text retrieval recall@1 results on Crossmodal-3600.
{NiceTabular}
l|l|ccc|cc|ccc|cc|c Winoground SugarCrepe-Add SugarCrepe-Replace SugarCrepe-Swap
Method Dataset Image Text Both Attribute Object Attribute Object Relation Attribute Object Avg.
CLIP [41] WIT-400M 10.8 25.0 7.5 66.8 78.5 81.1 93.4 66.9 64.6 60.0 55.5
DreamLIP [63] YFCC15M 14.7 26.2 9.7 78.3 80.3 81.3 91.0 72.9 77.5 66.9 59.9
FLAME YFCC15M 18.3 34.5 13.2 82.4 87.8 85.8 94.1 79.9 67.6 66.1 63.0
Table C3: Vision-language compositionality. FLAME demonstrates better fine-grained scene understanding capability.
{NiceTabular}
l|l|ccccc|ccccc Method N-Way K-shot 2-1 2-3 2-1 2-1 Avg. 5-1 5-3 5-1 5-1 Avg.
# Repetitions 0 0 1 3 0 0 1 3
Frozen [50] - 33.7 66.0 63.0 65.0 56.9 14.5 34.7 33.8 33.3 29.1
LQAE [35] GPT-3.5 35.2 68.2 68.5 68.7 60.2 15.7 35.9 31.9 36.4 30.0
V2L-Tokenizer [64] Llama-2-7B 76.3 91.2 84.0 84.4 84.0 44.8 91.8 73.9 82.2 73.2
SPAE [58] PaLM-2-340B 84.8 92.5 84.8 85.2 86.8 65.1 73.7 66.4 67.0 68.1
FLAME Mistral-7B 83.3 91.7 85.7 86.3 86.8 55.7 82.1 65.7 70.1 68.4
Table C4: Few-shot image comprehension on MiniImageNet. FLAME’s performance is either comparable to or surpasses that of SPAE.
C.1 Multilingual Image-to-Text Retrieval
We provide image-to-text retrieval results on Crossmodal-3600 [46] in Figure C1.
C.2 Vision-Language Compositionality
To evaluate the vision-language compositional understanding (i.e. whether the model understands the fine-grained atomic concepts that compose the scene), we conduct experiments on the Winoground [47] and SugarCrepe [18] benchmarks, with results presented in Table C. As shown, FLAME trained on YFCC15M significantly outperforms WIT-400M-trained CLIP [41] across all tasks, especially in relation understanding. Furthermore, when compared to the recent work [63] that utilizes multiple short captions, FLAME retains an advantage on 8 out of 10 tasks, achieving an average improvement of 3.1%. These results reveal that our multifaceted training promotes fine-grained semantic learning.
{NiceTabular}
c|c|cccccccccc|c|c Backbone Method Food-101 CIFAR-10 CIFAR-100 SUN397 Cars Aircraft DTD Pets Caltech-101 Flowers Average ImageNet
ViT-B/16 CLIP [41] 35.0 67.1 34.8 42.0 5.1 6.3 13.9 20.4 54.5 44.3 32.3 34.1
FLAME 61.8 86.1 56.7 66.8 10.7 10.3 54.9 40.7 78.9 51.7 51.9 51.5
ViT-L/14 CLIP [41] 42.2 66.8 33.1 45.3 2.7 2.3 19.7 26.4 65.1 53.1 35.7 36.3
FLAME
68.3 87.5 58.4 68.3 14.6 11.3 56.3 43.8 80.3 56.3 54.5 54.8
Table C5: Zero-shot classification on YFCC15M with ViT-L/14. FLAME consistently demonstrates substantial advantages over CLIP.
{NiceTabular}
c|c|ccccccc|c Backbone Method Food-101 CIFAR-10 CIFAR-100 Cars Aircraft DTD Caltech-101 Average
ViT-B/16 CLIP [41] 77.2 88.5 66.4 29.0 25.5 65.2 82.4 62.0
FLAME 85.9 95.0 81.0 54.3 39.3 76.8 92.5 75.0
ViT-L/14 CLIP [41] 74.4 88.9 69.7 27.3 26.1 63.7 86.4 62.3
FLAME
88.0 95.9 81.7 64.3 46.3 78.4 93.8 78.3
Table C6: Linear-probe classification on YFCC15M with ViT-L/14. FLAME benefits from the increased scale of the visual backbone.
C.3 Few-Shot Image Comprehension
Following SPAE [58] and V2L-Tokenizer [64], we conduct few-shot image comprehension experiments on both 2-way and 5-way MiniImageNet benchmarks. Our evaluation methodology involves converting image patches into words and then using the original LLM for in-context reasoning, without additional fine-tuning. Specifically, given a few-shot example image set $\{x_{i}\}_{i=1}^{S}$, we perform the vocabulary mapping process to obtain $\{C_{i}\}_{i=1}^{S}$, as mentioned in Section 4.4. By pairing each $C_{i}$ with its corresponding text answer $A_{i}$, we construct the in-context example $E=\{\langle C_{i},A_{i}\rangle\}_{i=1}^{S}$. We input this example and the new context $\tilde{C}$ to the LLM, which then outputs the answer $\tilde{A}$ for verification of correctness. This reasoning process can be formulated as $\tilde{A}=\mathrm{LLM}(E,\tilde{C})$. As shown in Table C, FLAME achieves average accuracies of 86.8% and 68.4% on 2-way and 5-way scenarios, respectively. These results are comparable to or surpass SPAE [58]. FLAME achieves these results using a smaller 7B backbone than SPAE’s 340B model.
C.4 Ablations on Individual Semantic Levels
{NiceTabular}
l|cc Long Avg.
Level I2T T2I
+Scene 50.5 47.4
+Interaction 52.9 50.3
+Entity 64.1 62.0
Table C7: Semantic decomposition.
To further clarify the reasonability of our semantic decomposition, we conduct ablations by orderly increasing individual semantic levels. The results are presented in Table C.4. Evidently, a systematic increase in semantic levels consistently yields positive outcomes. This also indicates that our rational prompt design is generalizable and does not lead to detrimental caption-prompt misalignment.
C.5 Preliminary Exploration of Scalability
To explore the scalability of FLAME, we increase the size of the visual encoder from the default ViT-B/16 to ViT-L/14 and train it on YFCC15M.
Table C.2 presents the zero-shot classification results, where FLAME continues to exhibit a substantial advantage over CLIP when using ViT-L/14, as evidenced by an improvement in ImageNet top-1 accuracy of 18.5% and an increase in downstream average accuracy of 18.8%. We also perform linear-probe classification experiments, with results shown in Table C.2.
{NiceTabular}
l|l|cccccc Evaluation
Method Training ar en it jp zh avg
CLIP [41] en 0.6 75.5 27.6 4.5 1.9 22.0
CN-CLIP [54] en+zh 0.1 32.5 8.0 16.3 53.4 22.1
NLLB-CLIP [51] en+multi 25.8 36.7 27.4 23.9 24.3 27.6
FLAME en 34.9 54.8 44.9 39.6 42.6 43.4
Table C8: Multilingual ImageNet1k classification with the “large" model variant. CLIP is trained on WIT-400M. CN-CLIP is pre-trained on WIT-400M and fine-tuned on 200M Chinese data. NLLB-CLIP is pre-trained on LAION-2B and fine-tuned on LAION-COCO-NLLB with 200 languages. FLAME is trained on YFCC15M.
In multilingual scenarios, the use of this larger backbone also proves beneficial, yielding a 2.8% improvement in average accuracy on the multilingual ImageNet1k classification benchmark. Detailed results and comparisons are provided in Table C.5.
These notable performance enhancements illustrate the scalability of FLAME, paving the way for future large-scale language-image pre-training.
D Visualizations
Figure D2 presents some visualizations of text-to-image retrieval on Urban-1k, comparing the results of FLAME with those of CLIP.
Figure D2: Visualizations of text-to-image retrieval on Urban-1k.